OpenStudy (anonymous):

If a coin is tossed 8 times what are the ODDS that it will be heads at least twice?

OpenStudy (anonymous):

PLEASE SOMEONE HELP ME! THIS IS A VERY IMPORTANT PROBLEM SET DUE TOMRW

OpenStudy (sidsiddhartha):

what are the ODDS that it will be heads at least twice this means what are the odds that it is heads once plus the odds that it is heads zero times.

OpenStudy (sidsiddhartha):

now heads 0 times means there will be only tails for 8 times so Heads 0 times would be 1/2^8

OpenStudy (anonymous):

i know the formula being the success over the failure. but i also know that the only way to know get heads at least twice would be to get it once or not at all. so how do i figure those numbers into an equation?

OpenStudy (sidsiddhartha):

and Heads 1 time would be (1/2^8) * (8)=(8/2^8) like this HTTTTTTT or THTTTTTT or TTHTTTTT, etc.

OpenStudy (sidsiddhartha):

now just Summing them up is 1/256 + 8/256 = 9/256

OpenStudy (sidsiddhartha):

do u get it so far?

OpenStudy (anonymous):

why did you multiply the fraction by 8?

OpenStudy (sidsiddhartha):

look Heads once 1/2, Tails 7 times 1/2^7 but it can go on like HTTTTTTT or THTTTTTT or TTHTTTTT or TTTHTTTT or TTTTHTTT or TTTTTHTT or TTTTTTHT or TTTTTTTH like this that is 8 times that is why u have to multiply 8

OpenStudy (anonymous):

oh ok that makes sense

OpenStudy (sidsiddhartha):

now ur ans is 1 minus the odds of 0 heads minus the odds of exactly 1 head. so \[prob=(1-\frac{ 9 }{ 256 })=247/256\]

OpenStudy (anonymous):

why is it 1?

OpenStudy (sidsiddhartha):

since we just calculated the odds of NOT meeting the objective

OpenStudy (anonymous):

thank you!

OpenStudy (sidsiddhartha):

no prob ^_^